### 27 Mar 2026 v1

# The adjoint state method for parametric definable optimization without smoothness or uniqueness

Abstract
Definable parametric optimization problems with possibly nonsmooth objectives, inequality constraints, and non-unique primal and dual solutions admit an adjoint state formula under a mere qualification condition. The adjoint construction yields a selection of a conservative field for the value function, providing a computable first-order object without requiring differentiation of the solution mapping. Through examples, we show that even in smooth problems, the formal adjoint construction fails without conservativity or definability, illustrating the relevance of these concepts to grasp theoretical aspects of the method. This work provides a tool which can be directly combined with existing primal-dual solvers for a wide range of parametric optimization problems.
Key words and phrases. Adjoint state method, conservative fields, nonsmooth optimization, o-minimal geometry, parametric optimization, variational analysis.
2020 Mathematics Subject Classification. Primary 90C31; Secondary 49J53, 03C64, 65K05, 90C26.
1 Introduction
Parametric optimization and the adjoint method.
Parametric optimization problems are central in optimization [9], control [31], and engineering [10]. They also arise naturally in modern machine learning, where one seeks to tune hyperparameters, architectures, regularization levels, or embedded optimization layers through the optimization of an induced value function; see [35, 4]. In these situations, the quantity of interest often has the parametric form
$f(\theta)\coloneq\inf_{x\in\mathbb{R}^{n}}\left\{F(x,\theta):G(x,\theta)\leq 0,H(x,\theta)=0\right\},$ (1.1)
where the parameter $\theta$ ranges in $\mathbb{R}^{q}$, $F\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$ is locally Lipschitz, the constraint functions $G\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{m}$ and $H\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{p}$ are continuously differentiable, and the inequality is understood coordinate-wise. A central difficulty in optimizing the value function $f$ lies in obtaining useful first-order information at a reasonable computational cost. In principle, if one denotes by
$x^{*}(\theta)\in\operatorname*{arg\,min}_{x\in\mathbb{R}^{n}}\left\{F(x,\theta):G(x,\theta)\leq 0,H(x,\theta)=0\right\},$
a selection of minimizers, the value writes $f(\theta)=F(x^{*}(\theta),\theta)$, thus, if the mapping $\theta\mapsto x^{*}(\theta)$ were differentiable, the chain rule would give
$\nabla f(\theta)=\nabla_{\!x}F(x^{*}(\theta),\theta)^{\!\mathsf{T}}\frac{dx^{*}(\theta)}{d\theta}+\nabla_{\!\theta}F(x^{*}(\theta),\theta).$
However, this approach raises both computational and conceptual difficulties. On the computational side, differentiating through a solver is generally expensive, while conceptually it relies on strong regularity assumptions, typically requiring an application of the implicit function theorem, which is often restrictive in practice.
The adjoint state method, introduced by Céa [11], provides a way to bypass the computation of the derivative of the solution mapping. It also avoids relying on the duality framework typically associated with convexity assumptions [37, 21]. Denoting $(x^{*}(\theta),\lambda^{*}(\theta),\mu^{*}(\theta))$ a solution of the KKT system of (1.1), the adjoint formula writes (see [11] for a formal derivation under equality constraints)
$\nabla f(\theta)=\nabla_{\!\theta}F(x^{*}(\theta),\theta)+\operatorname{Jac}_{\theta}\!G(x^{*}(\theta),\theta)^{\!\mathsf{T}}\lambda^{*}(\theta)+\operatorname{Jac}_{\theta}\!H(x^{*}(\theta),\theta)^{\!\mathsf{T}}\mu^{*}(\theta).$ (1.2)
Remarkably, at a negligible additional cost—see Remark 2.8()—associated with computing Lagrange multipliers, this formula does not involve the derivative of the primal-dual solution mapping. Yet, the theoretical validity of the analysis of Céa requires solutions of the KKT system to be unique and differentiable.
The adjoint formula in the definable world.
The objective of this work is to analyze the adjoint formula, under simple and versatile assumptions, that applies to a broad class of nonsmooth optimization problems with inequality constraints and possibly nonsmooth or non-unique primal and dual solutions. To this end, we show that the expression (1.2) carries strong variational information whenever the Mangasarian–Fromovitz constraint qualification holds and all the mappings $F,G,H$ belong to a common o-minimal structure (for example, are semialgebraic). We also discuss the computational cost of our approach, comparing it with less sophisticated methods and connecting it to automatic differentiation.
Our approach relies on two structural ingredients: definability and conservative calculus. Definability provides a natural framework that encompasses most optimization and machine learning models [8, 15], while offering powerful stability and regularity properties that are essential for our analysis. Conservative calculus, on the other hand, supplies generalized differential objects suited to nonsmooth calculus and has already led to strong applications in optimization and learning; see [8, 15, 34]. The use of conservative fields for the value function is essential here: the formal application of the adjoint formula in (1.2) produces artifacts, even for smooth data $F,G,H$ resulting in a smooth value function $f$, see the very simple Example 2.7(). This is very common in nonsmooth differential calculus. Nonetheless, the outputs of the adjoint formula (1.2) form a conservative field, in particular, artifacts only have a minor impact. We also point out that the natural connections between conservativity, differentiable programming, and algorithmic differentiation make our approach amenable to modern computational frameworks [7, 8].
Although primarily theoretical, this work provides a rigorous foundation for the practical use of the adjoint state method in settings where smoothness or uniqueness of solutions cannot be guaranteed. Such situations arise in a wide range of applications, including robust optimization [5], uncertainty quantification [1], optimal design of structures and systems [1], parametric decomposition [24, 20], and discretized neural ODEs [12]. A detailed investigation of these applications is beyond the scope of the present paper.
To conclude, we emphasize that the definable assumption is quite sharp. Indeed, in the absence of definability, the adjoint formula may fail to carry variational meaning for the value function, even for locally Lipschitz, path differentiable and (point-wise) differentiable objective function with smooth constraints. We provide an example showing that some structural assumption is required, justifying the o-minimal (definable) framework to ensure the validity of the adjoint formula. This framework is broad enough to cover most models that arise in applications. For example, most relevant neural networks are definable in some o-minimal structure [8, 15]. Beyond machine learning, many engineering and numerical models, when discretized, involve solely polynomials or definable functions.
Related work.
Let us briefly describe some related literature. In the continuous, infinite dimensional setting, extensions of the adjoint method to nonsmooth or non-unique solutions can be found in [16, 43]. They typically provide results on directional derivatives and rely on saddle-point-type arguments of a $t$-parametrized Lagrangian. The saddle point argument was removed later in [17, 18, 19]. A naive adaptation of these previous works to our discrete setting would yield, at best, only directional derivatives, rather than a directly computable gradient-like object for the value function, as provided by our approach. Moreover, it would require, when possible, solving costly optimization problems over the sets of primal and dual solutions. In nonlinear programming, related sensitivity results based on Lagrange multipliers likewise provide only upper and lower bounds on directional derivatives and entail similar optimization procedures over both primal and dual solution sets; see the work of Gauvin [23] and subsequent extensions [28, 22, 33, 3, 13, 38, 41]. Beyond directional derivatives, [27] and [36, Theorem 10.13] established that the Clarke subdifferential of the value function is included in a convex set constructed with the subdifferential of the Lagrangian; see the discussion in Section 4.2. However, as noted in Example 2.7(), this inclusion is strict in general, further confirming that subdifferentials are not the best paradigms to analyze the adjoint state method with multiple solutions. A limitation that this paper overcomes through conservative gradients.
Notations.
Let $\ell,L\in\mathbb{N}$ be generic natural numbers. We denote by $\langle{\cdot,\cdot}\rangle$ the canonical Euclidean scalar product on $\mathbb{R}^{\ell}$ and $\lVert{\cdot}\rVert$ its associated norm. For $r>0$ and $x\in\mathbb{R}^{\ell}$, we denote by $\mathcal{B}_{c}(x,r)$ the closed ball centered at $x$ with radius $r$, and by $\mathcal{B}(x,r)$ its open counterpart. Let $A\subset\mathbb{R}^{\ell}$. We write $\operatorname*{cl}A$, $\operatorname*{int}A$, and $\operatorname*{conv\!}A$ for the closure, interior, and convex hull of $A$, respectively. The Jacobian and partial Jacobian of a vector-valued function $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}^{L}$ at $z=(x,y)\in\mathbb{R}^{\ell}$ are denoted by $\operatorname{Jac}\varphi(z)$ and $\operatorname{Jac}_{x}\!\varphi(x,y)$, respectively. In the same way, the gradient and partial gradient of a real-valued function $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}$ at $z=(x,y)\in\mathbb{R}^{\ell}$ are written as $\nabla\varphi(z)$ and $\nabla_{\!x}\varphi(x,y)$, respectively. For $s,t\in\mathbb{R}$, we will sometimes use the notation $\varphi(t^{+})$ for $\lim_{s\downarrow t}\varphi(s)\coloneq\lim_{s>t,s\to t}\varphi(s)$ and the notation $\varphi(t^{-})$ for $\lim_{s\uparrow t}\varphi(s)\coloneq\lim_{s<t,s\to t}\varphi(s)$.
2 The adjoint state method: smooth case
This section studies the adjoint state method in the setting where the objective function $F$ is continuously differentiable, in the presence of equality and inequality constraints and possibly a continuum of minimizers. We focus here on the smooth case for simplicity and will expand on nonsmooth objectives in Section 4.
2.1 Preliminaries
We consider the following parametric nonlinear programming problem:
$\begin{aligned} \operatorname*{minimize}_{x\in\mathbb{R}^{n}}\quad&F(x,\theta)\\ \mathrm{subject\text{ }to}\quad&G(x,\theta)\leq 0,\\ &H(x,\theta)=0,\end{aligned}$ (2.1)
where $\theta\in\mathbb{R}^{q}$. The objective function $F:\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$ is locally Lipschitz, and the constraint mappings
$G(x,\theta)=\big(g_{1}(x,\theta),\ldots,g_{m}(x,\theta)\big)^{\!\mathsf{T}},\qquad H(x,\theta)=\big(h_{1}(x,\theta),\ldots,h_{p}(x,\theta)\big)^{\!\mathsf{T}}$
are continuously differentiable, the inequality being understood coordinate-wise. For any $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$, let $I(x,\theta)\subset\{1,\ldots,m\}$ be the set of active (inequality) constraints, i.e., for all $i\in I(x,\theta)$, $g_{i}(x,\theta)=0$. The Lagrangian of Problem (2.1), $\mathcal{L}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}$, writes
$\mathcal{L}(x,\lambda,\mu,\theta)\coloneq F(x,\theta)+\lambda^{\!\mathsf{T}}G(x,\theta)+\mu^{\!\mathsf{T}}H(x,\theta).$
Definition 2.1 (Mangasarian-Fromovitz constraint qualification).
Let $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $G(x,\theta)\leq 0,H(x,\theta)=0$. We say that the MFCQ holds at $(x,\theta)$, if the set
$\left\{\nabla_{\!x}h_{j}(x,\theta),j=1,\ldots,p\right\}$
is linearly independent and there exists a direction $d\in\mathbb{R}^{n}$ such that
$\nabla_{\!x}g_{i}(x,\theta)^{\!\mathsf{T}}d<0\text{ for all }i\in I(x,\theta)\text{ and }\operatorname{Jac}_{x}\!H(x,\theta)d=0.$
The MFCQ has an equivalent dual formulation. Indeed, the MFCQ is satisfied at $(x,\theta)$ if and only if
$\sum_{i\in I(x,\theta)}\lambda_{i}\nabla_{\!x}g_{i}(x,\theta)+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu=0,\text{ and }\lambda_{i}\geq 0,\forall\,i\in I(x,\theta)$
$\implies\mu=0,\lambda_{i}=0,\forall\,i\in I(x,\theta).$ (2.2)
Remark 2.2.
If the MFCQ is satisfied at $(x,\theta)$, then the following holds for the joint Jacobians in $x,\theta$,
$\operatorname{Jac}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}\!H(x,\theta)^{\!\mathsf{T}}\mu=0,\text{ and }\lambda_{i}\geq 0,\forall\,i\in I(x,\theta)$
$\implies\mu=0,\lambda_{i}=0,\forall\,i\in I(x,\theta).$ (2.3)
In other words, the MFCQ holds taking into account the joint dependence in $x,\theta$, not just the partial dependence in $x$
Assumptions 2.3.
The following assumptions hold throughout the current section, Section 2().
  1. (i)
    The objective function $F$ is continuously differentiable and is semialgebraic, i.e., its graph is described by polynomial inequalities and equalities. The constraint functions $G$ and $H$ are also semialgebraic.
  2. (ii)
    The MFCQ is satisfied at any $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $G(x,\theta)\leq 0,H(x,\theta)=0$.
  3. (iii)
    For all $\theta\in\mathbb{R}^{q}$, the solution set
    $\{x\in\mathbb{R}^{n}:F(x,\theta)=f(\theta),G(x,\theta)\leq 0,H(x,\theta)=0\}$
    is nonempty and bounded by $M_{\theta}\geq 0$. In addition, for all $\theta\in\mathbb{R}^{q}$, there exist a neighborhood $\mathcal{N}$ of $\theta$ and $M\geq 0$ such that for any $\bar{\theta}\in\mathcal{N}$, $M_{\bar{\theta}}\leq M$.
We consider a selection mapping, $\mathrm{solver_{p}}\colon\mathbb{R}^{q}\to\mathbb{R}^{n}$, such that for all $\theta\in\mathbb{R}^{q}$,
$\mathrm{solver_{p}}(\theta)\in\{x\in\mathbb{R}^{n}:F(x,\theta)=f(\theta),G(x,\theta)\leq 0,H(x,\theta)=0\},$ (2.4)
[…]
Let us define similarly a primal-dual solver, $\mathrm{solver_{pd}}\colon\mathbb{R}^{q}\to\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$, such that for all $\theta\in\mathbb{R}^{q}$,
$\mathrm{solver_{pd}}(\theta)$ $\in\big\{(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}:F(x,\theta)=f(\theta),G(x,\theta)\leq 0,H(x,\theta)=0,$
$\qquad\lambda\geq 0,G(x,\theta)^{\!\mathsf{T}}\lambda=0,\text{ and }$
$\qquad\nabla_{\!x}F(x,\theta)+\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu=0\big\}.$ (2.5)
[…]
2.2 The adjoint method
We propose the use of the adjoint formula from (1.2), without imposing any uniqueness requirement neither on the solutions to (2.1), nor on their associated dual/adjoint variables (Lagrange multipliers).
Algorithm ASM Adjoint State Method
[…]
3:Compute:
$u\leftarrow\nabla_{\!\theta}\mathcal{L}(x^{*},\lambda^{*},\mu^{*},\bar{\theta})=\nabla_{\!\theta}F(x^{*},\bar{\theta})+\operatorname{Jac}_{\theta}\!G(x^{*},\bar{\theta})^{\!\mathsf{T}}\lambda^{*}+\operatorname{Jac}_{\theta}\!H(x^{*},\bar{\theta})^{\!\mathsf{T}}\mu^{*}.$ (2.6)
4:$u$   (surrogate gradient for the value function)
Let us discuss the connection with the adjoint formula as originally formulated by Céa [11], under equality constraints. We have, for $(x^{*},\lambda^{*},\mu^{*})=\mathrm{solver_{pd}}(\bar{\theta})$, that
$\operatorname{Jac}_{x}\!G(x^{*},\bar{\theta})^{\!\mathsf{T}}\lambda^{*}+\operatorname{Jac}_{x}\!H(x^{*},\bar{\theta})^{\!\mathsf{T}}\mu^{*}=-\nabla_{\!x}F(x^{*},\bar{\theta}).$ (2.7)
In (2.7) and (2.6), one can recognize the operations of the adjoint state method proposed by [11, Equations (2.7) and (2.9)] with $m=0$ (no inequality constraint). The theoretical validity of the formula in [11], is based on the uniqueness and differentiability of both the primal and dual solutions, essentially requiring strong second order sufficiency conditions. The use of conservative fields enables us to consider the exact same formula as [11], without these strong conditions, allowing for possible nonsmooth or non-unique primal-dual solutions.
The following is our main result regarding the adjoint state method ASM() for smooth objectives. It is a particular case of Theorem 4.6() in Section 4(), where the result is proved.
Theorem 2.4.
Let Assumptions 2.3() stand. Then, the adjoint state method ASM() is a selection of a conservative field  for $f$, i.e.,
$\emph{ASM}(\theta)\in D_{f}(\theta)\quad\forall\,\theta\in\mathbb{R}^{q},$
where $D_{f}$ is a conservative field for $f$.
The conservative field $D_{f}$ in Theorem 2.4() is rigourously defined in Theorem 4.6(). Formally, it consists of all the possible outputs of the method ASM() for all possible choice of $\mathrm{solver_{pd}}$ as in the right hand side of (2.1).
Remark 2.5 (On conservative fields).
The result in Theorem 2.4() allows to invoke generic properties of conservative fields. Among them, let us highlight the following.
  1. (i)
    Conservative fields represent a notion of generalized gradients [8]. Theorem 2.4() provides a rigorous mathematical framework and analysis that justify the use of the adjoint state method without any uniqueness or differentiability assumptions on the primal/state and dual/adjoint solutions. The standard notions of subdifferential from nonsmooth analysis [36] are insufficient to describe the variational properties of the adjoint method under those conditions; see Example 2.7(). This justify the relevance of the notion of conservative field in this context.
  2. (ii)
    For general locally Lipschitz semialgebraic functions, like $f$, selections of conservative fields can be used for optimization. They carry sufficient variational information to serve as oracle for first-order methods with qualitative guaranties; see [8, 42]. This fact enables, for instance, parameter optimization for (2.1) through first-order methods by employing the adjoint state formula to efficiently compute a gradient-like oracle for $f$ at each iteration.
  3. (iii)
    Conservative fields are compatible with differentiable programming paradigms such as algorithmic differentiation [8, 7]. This means that, in practice, ASM() can be combined with further operations in a larger numerical program, in a way which is compatible with algorithmic differentiation as implemented for example in Python libraries such as Pytorch.
  4. (iv)
    The combination of [13, Proposition 2.1.2] and [8, Corollary 1] allows to obtain the following bound on Dini directional derivatives for any $\theta,d\in\mathbb{R}^{q}$:
    $\min_{v\in D_{f}(\theta)}\left\langle v,d\right\rangle\leq\liminf_{t\downarrow 0}\frac{f(\theta+td)-f(\theta)}{t}\leq\limsup_{t\downarrow 0}\frac{f(\theta+td)-f(\theta)}{t}\leq\max_{v\in D_{f}(\theta)}\left\langle v,d\right\rangle.$
    This is reminiscent of the sensitivity analysis performed, for example, in [23, 33], where more precise results on directional derivatives are proposed.
Remark 2.6 (On the semialgebraic assumption).
More generally, Theorem 2.4() and all its consequences in Remark 2.5() remain valid if all the functions, $F,G\text{ and }H$, are definable in the same o-minimal structure; see Definition 3.4(). This allows to include transcendental operations in the definition of $F,G\text{ and }H$, such as exponential or logarithm, and covers virtually all applications which can be considered in practice.
Example 2.7.
The following example shows that the (Clarke) subdifferential cannot capture all the information of the output of the adjoint formula in ASM() as the latter may produce artifacts which do not carry any standard variational meaning.
Set $n=2$ and $m=1$ with
$F(x,\theta)$ $=x_{1}x_{2},$ $G(x,\theta)=(x_{1},x_{2}-\theta,x_{2}+\theta)\!.^{\!\mathsf{T}}$
The constraints are expressed equivalently $x_{1}\leq 0$ and $x_{2}\leq-|\theta|$. We have $f(\theta)=0$ for all $\theta\in\mathbb{R}$, and the MFCQ holds. For $\bar{\theta}=0$, we have the following KKT solution
$x^{*}$ $=\begin{pmatrix}-1\\ 0\end{pmatrix}$, $\lambda^{*}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$,
such that $\nabla_{\!\theta}\mathcal{L}(x^{*},\lambda^{*},\bar{\theta})=1$, which is not a subgradient of $f$.
2.3 Comparison to alternative methods
Throughout this section $\bar{\theta}\in\mathbb{R}^{q}$ denotes a fixed set of parameters. We compare ASM() with the following two widely used approaches that are based on differentiating the solution to get the sensitivities of $f$. The first approach, which we will simply call automatic differentiation, uses reverse-mode algorithmic differentiation to obtain the derivative of the solution $x(\bar{\theta})$ by differentiating through $\mathrm{solver_{p}}$. The second is implicit differentiation, where the implicit function theorem is applied to the critical condition $\nabla\mathcal{L}_{1}(\mathrm{solver_{pd}}(\bar{\theta}))=0$, with $\mathcal{L}_{1}\colon(x,\lambda,\mu)\mapsto\mathcal{L}(x,\lambda,\mu,\bar{\theta})$ where $(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$ and $\bar{\theta}$ is fixed.
In terms of assumptions, it is immediately apparent that both methods require the uniqueness and differentiability of the solutions, while ASM() does not. For the purpose of comparing computational costs, let us consider the favorable settings ensuring the validity of both methods.
A straightforward way to implement ASM(), after getting a primal-dual solution from $\mathrm{solver_{pd}}$, is for the user to just compute $u=\texttt{autodiff}\mathcal{L}_{2}(\bar{\theta})$, with $\mathcal{L}_{2}\colon\theta\mapsto\mathcal{L}(\mathrm{solver_{pd}}(\bar{\theta}),\theta)$, where $\bar{\theta}$ is fixed, treated as a constant in the differentiation process. We use the notation Cost to denote the computational cost of implementing operations, which can be identified with computational time. According to the cheap gradient principle [26],
$\text{Cost}\big(\texttt{autodiff}\mathcal{L}_{2}(\bar{\theta})\big)\leq\omega\text{Cost}\big({\mathcal{L}_{2}}\big)=\omega\text{Cost}\big(\mathcal{L}\big),$
where $\omega>0$ is an absolute constant, $\text{Cost}\big(\texttt{autodiff}\mathcal{L}_{2}(\bar{\theta})\big)$ is the computational cost to calculate the derivative of $\mathcal{L}_{2}$ through reverse-mode automatic differentiation, and $\text{Cost}\big(\mathcal{L}_{2}\big)$ and $\text{Cost}\big(\mathcal{L}\big)$ the computational cost to evaluate the value of the functions $\mathcal{L}_{2}$ and $\mathcal{L}$, respectively. Therefore, the total cost of the adjoint state method ASM() is upper bounded as follows:
$\text{Cost}\big(\text{ASM}\big)\leq\text{Cost}\big(\mathrm{solver_{pd}}\big)+\omega\text{Cost}\big(\mathcal{L}\big).$ (2.8)
Typically, the term $\text{Cost}\big(\mathrm{solver_{pd}}\big)$ dominates on the right hand side.
Remark 2.8.
In most nonlinear programming solvers, such as SNOPT [25] and IPOPT [40], dual variables are already computed internally as part of satisfying the KKT optimality conditions, or because of the primal-dual nature of the underlying algorithms. Consequently, in such a situation, getting the dual solution in addition to the primal solution, requires little to no additional computational effort, $\text{Cost}\big(\mathrm{solver_{pd}}\big)$ is the same as $\text{Cost}\big(\mathrm{solver_{p}}\big)$ and the adjoint state formula ASM is readily implementable with minimal code extension.
Both vanilla automatic differentiation and implicit differentiation are respectively based on the following chain rules,
$\nabla f(\bar{\theta})$ $=\nabla_{\!x}F\big(\mathrm{solver_{p}}(\bar{\theta}),\bar{\theta}\big)^{\!\mathsf{T}}\operatorname{Jac}\big(\mathrm{solver_{p}}(\bar{\theta})\big)+\nabla_{\!\theta}F\big(\mathrm{solver_{p}}(\bar{\theta}),\bar{\theta}\big)$
$=\nabla\mathcal{L}_{1}\big(\mathrm{solver_{pd}}(\bar{\theta})\big)^{\!\mathsf{T}}\operatorname{Jac}\big(\mathrm{solver_{pd}}(\bar{\theta})\big)+\nabla\mathcal{L}_{2}(\bar{\theta}),$
The computational cost of automatic differentiation can be roughly bounded as follows:
$\text{Cost}\big(\text{AD}\big)$ $=\text{Cost}\big(\mathrm{solver_{p}}\big)+\text{Cost}\big(\texttt{autodiff}F(\mathrm{solver_{p}}(\bar{\theta}),\bar{\theta})\big)+\text{Cost}\big(\texttt{autodiff}\,\mathrm{solver_{p}}(\bar{\theta})\big)$
$\leq\text{Cost}\big(\mathrm{solver_{p}}\big)+\omega\text{Cost}\big(F\big)+\omega\text{Cost}\big(\mathrm{solver_{p}}\big).$ (2.9)
Again, in typical situations the terms $\text{Cost}\big(\mathrm{solver_{p}}\big)$ dominates. Finally, assuming all regularity conditions for its application [2], implicit differentiation has the following approximate bound on its computational cost:
$\text{Cost}\big(\text{ID}\big)$ $=\text{Cost}\big(\mathrm{solver_{pd}}\big)+\text{Cost}\big(\texttt{autodiff}\mathcal{L}(\mathrm{solver_{pd}}(\bar{\theta}),\bar{\theta})\big)$
$\qquad+\text{Cost}\Big(\texttt{autodiff}\big(\texttt{autodiff}\mathcal{L}\big)(\mathrm{solver_{pd}}(\bar{\theta}),\bar{\theta})\Big)+\text{Cost}\big(\text{LE}\big)$
$\leq\text{Cost}\big(\mathrm{solver_{pd}}\big)+\text{Cost}\big(\text{LE}\big)+\omega(n+q+1)\text{Cost}\big(\mathcal{L}\big),$ (2.10)
where Cost(LE) is the computational cost of inverting a linear system.
The choice of automatic differentiation and implicit differentiations in (2.3) and (2.3) depends on the number of variables and number of constraints, and possibly further characteristics of the problem. Depending on the setting, one solution may be preferable to the other, both in terms of worst case upper bounds and in practice. Note that this is a high level analysis and the actual performance may depend on additional computational factors, such as how second order derivatives are computed, how linear equations are solved, and potential approximation and parallelization strategies. Nonetheless this provides a reasonable upper bound on the computational burden for both automatic differentiation and implicit differentiation. Let us also mention that automatic differentiation requires to have access to the code of $\mathrm{solver_{p}}$, which may not be available for closed-source software.
Comparing with Equations (2.8), it is obvious that the worst case computational cost of ASM() is lower, and possibly much lower, than either automatic differentiation or implicit differentiation, making it a strong competitor for the purpose of differentiating the value function of nonlinear programs.
3 Elements of nonsmooth analysis and o-minimality
In full generality, our study is conducted in a setting where the objective function $F$ is merely locally Lipschitz. Accordingly, we begin by introducing tools required to handle nonsmoothness.
3.1 Conservative set-valued fields
A set-valued map or multifunction $D\colon\mathbb{R}^{\ell}\rightrightarrows\mathbb{R}^{s}$ is a mapping from $\mathbb{R}^{\ell}$ to the set of all subsets of $\mathbb{R}^{s}$. The graph of $D$ is given by $\operatorname*{gph}D:=\{(x,y)\in\mathbb{R}^{\ell}\times\mathbb{R}^{s}:y\in D(x)\}$. $D$ is locally bounded at $x\in\mathbb{R}^{\ell}$ if there exist a neighborhood $\mathcal{N}$ of $x$ and $r>0$ such that $\bigcup_{z\in\mathcal{N}}D(z)\subset\mathcal{B}_{c}(0,r)$. $D$ is graph closed if $\operatorname*{gph}D$ is a closed subset of $\mathbb{R}^{\ell}\times\mathbb{R}^{s}$. Equivalently, $D$ is graph closed if for all $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{\ell}$ and all $(y_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{s}$ such that $x_{k}\xrightarrow[k\to\infty]{}x$, $y_{k}\xrightarrow[k\to\infty]{}y$ and $y_{k}\in D(x_{k})$ for any $k\in\mathbb{N}$, it follows that $y\in D(x)$.
Definition 3.1 (Clarke generalized gradients [13]).
Let $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}$ be a locally Lipschitz function. By Rademacher’s theorem, $\varphi$ is differentiable on a full measure subset of $\mathbb{R}^{\ell}$, say $\Omega$. Then, the Clarke subdifferential of $\varphi$ is the set-valued map $\partial^{c}\varphi\colon\mathbb{R}^{\ell}\rightrightarrows\mathbb{R}^{\ell}$ defined as
$x\mapsto\operatorname*{conv\!}\,\{u\in\mathbb{R}^{\ell}:\exists\,(u_{k})_{k\in\mathbb{N}}\subset\Omega,u_{k}\xrightarrow[k\to\infty]{}x\text{ and }\nabla\varphi(u_{k})\xrightarrow[k\to\infty]{}u\}.$
The Clarke subdifferential of $\varphi$, like for any other locally Lipschitz real-valued function, has nonempty, convex and compact values, and is closed graph and locally bounded.
Given $(y,z)\in\mathbb{R}^{\ell}$, the Clarke partial subdifferential $\partial^{c}_{y}\varphi(y,z)$ of $\varphi$ at $(y,z)$ with respect to $y$ is the Clarke subdifferential of the function $\varphi(\cdot,z)$ at $y$.
Our analysis revolves around the notion of conservative gradients that we present in the next definition. Let $I\subset\mathbb{R}$ be a nonempty interval and $u\colon I\to\mathbb{R}^{\ell}$ be a curve. The curve $u$ is said to be absolutely continuous if $u$ is differentiable Lebesgue-almost everywhere on $I$, its derivative $\dot{u}$ is Lebesgue integrable on $I$, and $u(b)-u(a)=\int_{a}^{b}\dot{u}(t)dt$, for all $a,b\in I$.
Definition 3.2 (Conservative fields [8]).
Let $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}$ be a locally Lipschitz function and $D_{\varphi}\colon\mathbb{R}^{\ell}\to\mathbb{R}^{\ell}$ be a locally bounded and graph closed set-valued map with nonempty values. We say that $D_{\varphi}$ is a conservative field for $\varphi$ if for any absolutely continuous curve $u\colon[0,1]\to\mathbb{R}^{\ell}$,
$\frac{d}{dt}\varphi(u(t))=\langle{\dot{u}(t),v}\rangle\quad\forall\,v\in D_{\varphi}(u(t)),$ (3.1)
for almost all $t\in[0,1]$. Any such $\varphi$ is called path differentiable.
3.2 Definition of o-minimality
We briefly recall the concepts of tame geometry relevant to the class of mappings (definable mappings) considered in the present paper. Important works on this topic include [39, 14].
Definition 3.3 (o-minimal structures).
An o-minimal structure on $(\mathbb{R},+,\cdot)$ is a collection of families $\mathcal{O}=(\mathcal{O}_{\ell})_{\ell\in\mathbb{N}}$, where for each $\ell\in\mathbb{N}$, the set $\mathcal{O}_{\ell}$ is a family of subsets of $\mathbb{R}^{\ell}$, satisfying the following properties:
  1. (i)
    $\mathcal{O}_{\ell}$ is closed under complementation, finite unions, and finite intersections.
  2. (ii)
    If $A\in\mathcal{O}_{\ell}$, then both $A\times\mathbb{R}$ and $\mathbb{R}\times A$ belong to $\mathcal{O}_{\ell+1}$.
  3. (iii)
    Let $\pi:\mathbb{R}^{\ell+1}\to\mathbb{R}^{\ell}$ be the canonical projection. If $A\in\mathcal{O}_{\ell+1}$, then $\pi(A)\in\mathcal{O}_{\ell}$.
  4. (iv)
    The family $\mathcal{O}_{\ell}$ contains all real algebraic subsets of $\mathbb{R}^{\ell}$, which are sets of the form
    $\{x\in\mathbb{R}^{\ell}:\phi(x)=0\},$
    where $\phi\colon\mathbb{R}^{\ell}\to\mathbb{R}$ is a polynomial function.
  5. (v)
    The sets in $\mathcal{O}_{1}$ are precisely the finite unions of points and intervals.
Definition 3.4 (Definable sets and definable mappings).
A subset of $\mathbb{R}^{\ell}$ that belongs to an o-minimal structure $\mathcal{O}$ is said to be definable in $\mathcal{O}$. A function or a set-valued map is definable in $\mathcal{O}$ if its graph is definable in $\mathcal{O}$.
Example 3.5.
The simplest o-minimal structure is given by the class of real semialgebraic sets. Recall that a set $A\subset\mathbb{R}^{\ell}$ is called semialgebraic if it is a finite union of sets of the form
$\bigcap_{k=1}^{K}\{x\in\mathbb{R}^{\ell}:\phi_{k}(x)<0,\;\varphi_{k}(x)=0\},$
where $K\geq 1$ and $\phi_{k}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ and $\varphi_{k}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ are polynomial functions for all $k\in\{1,\ldots,K\}$. Therefore, the semialgebraic functions and semialgebraic set-valued maps are functions and set-valued maps, respectively, whose graphs are semialgebraic sets.
From now on, we fix an o-minimal structure $\mathcal{O}$. And unless stated otherwise, all definable sets and definable mappings are understood to be definable in $\mathcal{O}$.
4 The adjoint state method: general conservative case
4.1 Preliminaries
Define the set-valued map $R:\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}$ of feasible point of Problem (2.1) as
$R(\theta)\coloneq\left\{x\in\mathbb{R}^{n}:G(x,\theta)\leq 0,H(x,\theta)=0\right\},$
and the set-valued map $\mathrm{Sol_{p}}\colon\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}$ of primal solutions of (2.1) as
$\mathrm{Sol_{p}}(\theta)\coloneq\operatorname*{arg\,min}\left\{F(x,\theta):x\in R(\theta)\right\}.$
Remark 4.1.
By continuity of $G$ and $H$, the set-valued map $R$ is graph closed.
Our main construction is based on the following.
Assumptions 4.2 (Standing assumptions).
  1. (i)
    The objective function $F$ is definable with definable conservative field $D_{F}$. The constraint functions $G$ and $H$ are also definable. In that case, we call Problem (2.1) a parametric definable program.
  2. (ii)
    The MFCQ is satisfied at any $(\theta,x)\in\operatorname*{gph}R$, i.e., the MFCQ holds throughout the feasible set $R(\theta)$ for any $\theta\in\mathbb{R}^{d}$.
  3. (iii)
    The set-valued map $\mathrm{Sol_{p}}$ has nonempty values and is locally bounded.
Remark 4.3.
  1. (i)
[…]
  2. (ii)
[…]
We set $D_{F(\cdot,\theta)}\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$, the projection on the $x$ component of $D_{F}$, which provides a conservative for the partial function $F(\cdot,\theta)$, for fixed $\theta\in\mathbb{R}^{q}$.
Let $\operatorname*{Mult}\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{p}$ be the set-valued map providing Lagrange multipliers satisfying stationarity and complementarity from the KKT system
$(x,\theta)$ $\mapsto\Big\{(\lambda,\mu)\in\mathbb{R}^{m}\times\mathbb{R}^{p}:0\in\operatorname*{conv\!}D_{F(\cdot,\theta)}(x)+\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,$
$\qquad\quad\lambda\geq 0,\lambda^{\!\mathsf{T}}G(x,\theta)=0\Big\},$
We then define the set-valued map of primal-dual solutions $\mathrm{Sol_{pd}}\colon\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$ as
$\mathrm{Sol_{pd}}(\theta)\coloneq\{(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}:x\in\mathrm{Sol_{p}}(\theta),(\lambda,\mu)\in\operatorname*{Mult}(x,\theta)\}.$
Proposition 4.4.
The set-valued map of primal-dual solutions $\mathrm{Sol_{pd}}$ has nonempty values and is locally bounded.
For all $\theta\in\mathbb{R}^{q}$ and for every $x\in\mathrm{Sol_{p}}(\theta)$, we have the conservative KKT conditions given by [29, Theorem 3.3] and [8, Corollary 1]. Then $\operatorname*{Mult}(\theta,x)\neq\varnothing$ and $\mathrm{Sol_{pd}}(\theta)$ is nonempty.
Let $\theta\in\mathbb{R}^{q}$. For the sake of contradiction, let’s assume that $\mathrm{Sol_{pd}}$ is not locally bounded around $\theta$. It follows that there exist $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ and $(x_{k},\lambda_{k},\mu_{k})\in\mathrm{Sol_{pd}}(\theta_{k})$, for all $k\in\mathbb{N}$, such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta$ and $\lVert{(x_{k},\lambda_{k},\mu_{k})}\rVert\xrightarrow[k\to\infty$$\infty$. Since $\mathrm{Sol_{p}}$ is locally bounded, we have that $(x_{k})_{k\in\mathbb{N}}$ is bounded and $\lVert{(\lambda_{k},\mu_{k})}\rVert\xrightarrow[k\to\infty$$\infty$. Passing to subsequences if necessary, we can assume that $\lVert{(\lambda_{k},\mu_{k})}\rVert\neq 0$ for all $k\in\mathbb{N}$, and $x_{k}\xrightarrow[k\to\infty]{}x$. By Lemma A.2(), we get $x\in\mathrm{Sol_{p}}(\theta)$. Let, for all $k\in\mathbb{N}$, $\alpha_{k}=\frac{\lambda_{k}}{\lVert{(\lambda_{k},\mu_{k})}\rVert}$ and $\beta_{k}=\frac{\mu_{k}}{\lVert{(\lambda_{k},\mu_{k})}\rVert}$. So for any $k\in\mathbb{N}$, $\lVert{\alpha_{k}}\rVert\leq 1$ and $\lVert{\beta_{k}}\rVert\leq 1$, and, without loss of generality, we can also assume $\alpha_{k}\xrightarrow[k\to\infty]{}\alpha$ and $\beta_{k}\xrightarrow[k\to\infty]{}\beta$. For all $k\in\mathbb{N}$, we have that $\lVert{(\alpha_{k},\beta_{k})}\rVert=1$. It is clear that $\lVert{(\alpha,\beta)}\rVert=1$. By definition of the set-valued map $\operatorname*{Mult}$, we know that, for every $k\in\mathbb{N}$,
$\exists\,v_{k}\in\operatorname*{conv\!}D_{F(\cdot,\theta_{k})}(x_{k}),\operatorname{Jac}_{x}\!G(x_{k},\theta_{k})^{\!\mathsf{T}}\alpha_{k}+\operatorname{Jac}_{x}\!H(x_{k},\theta_{k})^{\!\mathsf{T}}\beta_{k}=-\frac{v_{k}}{\lVert{(\alpha_{k},\beta_{k})}\rVert}.$ (4.1)
Using the fact that $(x,\theta)\mapsto D_{F(\cdot,\theta)}(x)$ is locally bounded (because $D_{F}$ is locally bounded), we have that $(v_{k})_{k\in\mathbb{N}}$ is bounded. Again, passing to a subsequence if needed, we have that $v_{k}\xrightarrow[k\to\infty]{}v$. Taking the limit in (4.1), we can write
$\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\alpha+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\beta=0.$ (4.2)
Furthermore, for all $k\in\mathbb{N}$ and $i\in\{1,\ldots,m\}$,
$\alpha_{k,i}\geq 0\;\text{ and }\;\alpha_{k,i}g_{i}(x_{k},\theta_{k})=0.$
Passing to the limit, we get, for all $i\in\{1,\ldots,m\}$,
$\alpha_{i}\geq 0\;\text{ and }\;\alpha_{i}g_{i}(x,\theta)=0.$ (4.3)
Given Equations (4.2) and (4.3), the dual form of the MFCQ in (2.1) imposes $\alpha=\beta=0$. This contradicts the fact that $\lVert{(\alpha,\beta)}\rVert=1$. Thus $\mathrm{Sol_{pd}}$ is locally bounded around $\theta$. ∎
Remark 4.5.
For any $\theta\in\mathbb{R}^{q}$, thanks to Proposition 4.4(), $\mathrm{Sol_{pd}}(\theta)\neq\varnothing$ and we have
$f(\theta)=F(\mathrm{Sol_{p}}(\theta),\theta)=\mathcal{L}(\mathrm{Sol_{pd}}(\theta),\theta).$ (4.4)
The last equality in (4.4) is due to the fact that, for all $(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta)$, one has by complementarity
$x\in\mathrm{Sol_{p}}(\theta),\;\lambda^{\!\mathsf{T}}G(x,\theta)=0,\text{ and }H(x,\theta)=0.$
Thus, $F(\mathrm{Sol_{p}}(\theta),\theta)=F(x,\theta)=\mathcal{L}(x,\lambda,\theta)=\mathcal{L}(\mathrm{Sol_{pd}}(\theta),\theta)$.
4.2 Parametric subdifferentiation of the value function
As discussed earlier, [36, Theorem 10.13] provides a formula for the subdifferential of $f$, which has a strong connection with the adjoint formula in (1.2). Let us make this explicit in the context of Problem (1.1). We will consider, only in this section, the extended formulation as follows
$\bar{F}\colon\mathbb{R}^{n}\times\mathbb{R}^{q}$ $\quad\to\quad\mathbb{R}\cup\{\infty\}$
$(x,\theta)$ $\quad\mapsto\quad F(x,\theta)+\delta_{\mathbb{R}^{m}_{-}}(G(x,\theta))+\delta_{\{0^{p}\}}(H(x,\theta)),$
where, for $C\in\mathbb{R}^{\ell}$ with $\ell\in\mathbb{N}$, $\delta$ is the indicator function of $C$, i.e., $\delta_{C}(y)=0$ if $y\in C$ and $\delta_{C}(y)=\infty$ otherwise, and $\{0^{p}\}$ denotes the origin in $\mathbb{R}^{p}$. Problem (1.1) is equivalent to unconstrained partial minimization of $\bar{F}$ so that this fits the context of [36, Theorem 10.13].
Under Assumption 4.2 the constraint qualification holds jointly in $x,\theta$, see Remark 2.2 and the chain rule [36, Theorem 10.6] applies to the indicators in the definition of $\bar{F}$. Combining this with the sum rule [36, Exercise 10.10], we have, for all $(x,\theta)\in\operatorname*{gph}R$,
$\partial^{c}\bar{F}(x,\theta)$ $\subset\partial^{c}F(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta),\operatorname{Jac}_{\theta}\!G(x,\theta)\right)^{\!\mathsf{T}}N_{\mathbb{R}^{m}_{-}}(G(x,\theta))$
$\qquad+\left(\operatorname{Jac}_{x}\!H(x,\theta),\operatorname{Jac}_{\theta}\!H(x,\theta)\right)^{\!\mathsf{T}}N_{\{0^{p}\}}(H(x,\theta)),$
where, for $C\subset\mathbb{R}^{\ell}$ convex with $\ell\in\mathbb{N}$ and $y\in\mathbb{R}^{\ell}$, $N_{C}(y)$ is the normal cone of $C$ at $y$. In our case the elements of the normal cones are given by the Lagrange multipliers $\lambda,\mu$ in $\mathrm{Sol_{pd}}(\theta)$. Putting everything together, the parametric optimality condition, [36, Theorem 10.13], ensures that, under Assumptions 4.2(), for all $\theta\in\mathbb{R}^{q}$,
$\partial^{c}f(\theta)\subset D_{f}(\theta)\coloneq\operatorname*{conv\!}\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta),$ (4.5)
$(0,u)\in\partial^{c}F(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right)\Big\}.$
Importantly, for all $\theta\in\mathbb{R}^{q}$ and for all $x\in\mathrm{Sol_{p}}(\theta)$, thanks to the parametric Fermat’s rule [36, Example 10.12], there exists $u\in\mathbb{R}^{q}$ such that $(0,u)\in\partial^{c}\bar{F}(x,\theta)$; hence $D_{f}(\theta)$ is nonempty.
However, as already stated earlier, the inclusion in (4.5) is not sharp in general and the generalized adjoint formula produces artifacts, see Example 2.7. We show in the coming section that, under Assumptions 4.2(), the set-valued map $D_{f}$ is a conservative field for $f$, providing a conservative adjoint formula.
4.3 Conservative adjoint formula
This is our main result. We consider a slightly more general version of $D_{f}$ in (4.5), allowing for a general conservative field for $F$, beyond $\partial^{c}F$. The motivation for the proposed form of $D_{f}$ follows directly from the preceeding discussion in Section 4.2.
Theorem 4.6 (Main result).
It is now assumed that $F$ is a general locally Lipschitz function, not necessary differentiable. Recall that Assumptions 4.2() hold, that $D_{F}$ is a conservative field for $F$ and that the value function $f\colon\mathbb{R}^{q}\to\mathbb{R}$ is equal to
$\theta\mapsto F(\mathrm{Sol_{p}}(\theta),\theta).$
Define $D_{f}:\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{q}$ by
$\theta\mapsto\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta),$
$(0,u)\in\operatorname*{conv\!}D_{F}(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right)\Big\}.$
Then $D_{f}$ is a conservative field for $f$, and in particular, $\partial^{c}f(\theta)\subset\operatorname*{conv\!}D_{f}(\theta).$
Example 4.7.
The Clarke subdifferential is known to be conservative for definable functions [15]. Therefore, we can take the Clarke subdifferential of $F$ as the conservative $D_{F}$ for $F$ in Theorem 4.6() and everywhere $D_{F}$ is used. Optionally, for any $\theta\in\mathbb{R}^{q}$, we can replace $D_{F(\cdot,\theta)}$ by the Clarke partial subdifferential $\partial^{c}_{x}F(\cdot,\theta)$, which is contained in the projection of the joint sudifferential [13, Proposition 2.3.16].
The proof follows similar lines as [34, Theorem 12]. Define the set-valued map $D_{\mathcal{L}}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}\times\mathbb{R}^{q}$ as
$D_{\mathcal{L}}(y,\theta)$ $=D_{\mathcal{L}}(x,\lambda,\mu,\theta)$
$\coloneq D_{F}(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right),$
where we use the shorthand notation $y=(x,\lambda,\mu)$. Note that $D_{\mathcal{L}}$ is a partial conservative field for $\mathcal{L}$, which ignores the dependency in $\lambda,\mu$. This will be possible thanks to complementarity and the structure of $\mathcal{L}$.
As $D_{\mathcal{L}}$, $\mathrm{Sol_{p}}$ and $\mathrm{Sol_{pd}}$ can be described by first-order formulas, they are definable; see [14, Theorem 1.13]. By definition, the definable conservative field $D_{F}$ is graph closed and compact valued and locally bounded. By the parametric Fermat’s rule in [36, Example 10.12], $D_{\mathcal{L}}$ has nonempty values. It is also immediate that $D_{\mathcal{L}}$ is definable, graph closed, locally bounded, and compact valued by Proposition 4.4.
According to [34, Theorem 2 and Lemma 6], to finish the proof, it remains to show that $D_{f}$ has the chain rule property along continuously differentiable and definable curves. So let $t\mapsto\theta(t)$ be a continuously differentiable and definable curve in $\mathbb{R}^{q}$. The following definable selections are possible using [39, paragraph 4.5].
We fix for now $t\mapsto y(t)\coloneq(x(t),\lambda(t),\mu(t))$ a definable selection in $t\rightrightarrows\mathrm{Sol_{pd}}(\theta(t))$, with $t\mapsto x(t)$ a definable selection in $t\rightrightarrows\mathrm{Sol_{p}}(\theta(t))$ and $t\mapsto(\lambda(t),\mu(t))$ a definable selection in $t\rightrightarrows\operatorname*{Mult}(x(t),\theta(t))$. Let $t\mapsto T(t)\coloneq(v(t),u(t))$ be a definable selection in $t\rightrightarrows D_{\mathcal{L}}(y(t),\theta(t))$. There is a definable selection $t\mapsto(z(t),w(t))$ in $t\rightrightarrows D_{F}(x(t),\theta(t))$ such that for all $t$
$v(t)$ $=z(t)+\operatorname{Jac}_{x}\!G(x(t),\theta(t))^{\!\mathsf{T}}\lambda(t)+\operatorname{Jac}_{x}\!H(x(t),\theta(t))^{\!\mathsf{T}}\mu(t)$
$u(t)$ $=w(t)+\operatorname{Jac}_{\theta}\!G(x(t),\theta(t))^{\!\mathsf{T}}\lambda(t)+\operatorname{Jac}_{\theta}\!H(x(t),\theta(t))^{\!\mathsf{T}}\mu(t).$
Since definable curves are piecewise continuously differentiable, all considered selections as well as
$t\mapsto f(\theta(t))=F(x(t),\theta(t))=\mathcal{L}(x(t),\lambda(t),\theta(t)),$
are differentiable everywhere except at finitely many points, say at $t_{1},t_{2},\ldots,t_{M}$. By Proposition 4.4(), $\mathrm{Sol_{pd}}$ is locally bounded. It follows that $t\mapsto y(t)=(x(t),\lambda(t),\mu(t))$ has left and right limits everywhere and can be extended to an absolute continuous curve on $t_{\ell-1},t_{\ell}]$, for any $\ell\in\{2,\dots,M\}$.
Using the equalities in (4.4), the continuity of $f$ given by Lemma A.3(), and the continuity of $F$, $G$ and $H$, we have
$f(\theta(t_{\ell}))-f(\theta(t_{\ell-1}))$ $=\lim_{t\uparrow t_{\ell}}\mathcal{L}(x(t),\lambda(t),\theta(t))-\lim_{t\downarrow t_{\ell-1}}\mathcal{L}(x(t),\lambda(t),\theta(t))$
$=F(x(t_{\ell}^{-}),\theta(t_{\ell}))-F(x(t_{\ell-1}^{+}),\theta(t_{\ell-1}))$
$\qquad+\lambda(t_{\ell}^{-})^{\!\mathsf{T}}G(x(t_{\ell}^{-}),\theta(t_{\ell}))+\mu(t_{\ell}^{-})^{\!\mathsf{T}}H(x(t_{\ell}^{-}),\theta(t_{\ell}))$
$\qquad-\lambda(t_{\ell-1}^{+})^{\!\mathsf{T}}G(x(t_{\ell-1}^{+}),\theta(t_{\ell-1}))-\mu(t_{\ell-1}^{+})^{\!\mathsf{T}}H(x(t_{\ell-1}^{+}),\theta(t_{\ell-1}))$
$=\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}F(x(t),\theta(t))\,dt+\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}\left[\lambda(t)^{\!\mathsf{T}}G(x(t),\theta(t))\right]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}\left[\mu(t)^{\!\mathsf{T}}H(x(t),\theta(t))\right]dt.$
Using the fact that $D_{F}$ is conservative for $F$ in the first integral and differentiating in the others give
$f(\theta(t_{\ell}))-f(\theta(t_{\ell-1}))=\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\dot{\theta}(t),w(t)}\rangle+\langle{\dot{x}(t),z(t)}\rangle\big]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\dot{\lambda}(t),G(x(t),\theta(t))}\rangle+\langle{\dot{\mu}(t),H(x(t),\theta(t))}\rangle\big]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\lambda(t),\operatorname{Jac}_{\theta}\!G(x(t),\theta(t))\dot{\theta}(t)}\rangle+\langle{\mu(t),\operatorname{Jac}_{\theta}\!H(x(t),\theta(t))\dot{\theta}(t)}\rangle\big]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\lambda(t),\operatorname{Jac}_{x}\!G(x(t),\theta(t))\dot{x}(t)}\rangle+\langle{\mu(t),\operatorname{Jac}_{x}\!H(x(t),\theta(t))\dot{x}(t)}\rangle\big]dt.$ (4.6)
We have that, for any $\ell\in\{2,\dots,M\}$, on $(t_{\ell-1},t_{\ell})$,
$t\mapsto\langle{\dot{\mu}(t),H(x(t),\theta(t)))}\rangle\equiv t\mapsto\langle{\dot{\lambda}(t),G(x(t),\theta(t))}\rangle\equiv 0.$ (4.7)
This follows from the fact that, for any $t\in(t_{\ell-1},t_{\ell})$, all components of $G(x(t),\theta(t))$ are zero except those corresponding to inactive constraints $g_{i}$, with $i\in\{1,\ldots,m\}\setminus I(x(t),\theta(t))$. Moreover, the corresponding components $\dot{\lambda}_{i}(t)$ of $\dot{\lambda}(t)$ are also zero. Indeed, by continuity of $G$, for every $s$ in some open interval $\mathcal{I}$ containing $t$, $g_{i}(\theta(s),x(s))<0$ and $\lambda_{i}(s)=0$ for any $i$ such that $g_{i}(\theta(t),x(t))<0$. Hence, $\dot{\lambda}_{i}(t)=0$ for all such $i$.
Using (4.7), removing the finitely many discontinuous points and putting everything together, Equation (4.6) gives, for almost every $t$,
$\frac{d}{dt}f(\theta(t))=\langle{\dot{x}(t),v(t)}\rangle+\langle{\dot{\theta}(t),u(t)}\rangle.$ (4.8)
Since $D_{\mathcal{L}}$ is definable and compact valued, Lemma A.4() says that there is a sequence of definable selections that is dense in $D_{\mathcal{L}}(y(t),\theta(t))$, for all $t$. Each element of the sequence satisfies (4.8) for almost all $t$. Combining the closure in Lemma A.4(), with the fact that a countable union of Lebesgue negligible set is Lebesgue negligible, we obtain, for almost every $t$,
$\frac{d}{dt}f(\theta(t))=\langle{\dot{x}(t),v}\rangle+\langle{\dot{\theta}(t),u}\rangle\quad\forall\,(v,u)\in D_{\mathcal{L}}(y(t),\theta(t)).$
Then, we have, for almost every $t$,
$\frac{d}{dt}f(\theta(t))=\langle{\dot{\theta}(t),u}\rangle\quad\forall\,u\in\mathbb{R}^{q}\text{ such that }(0,u)\in D_{\mathcal{L}}(y(t),\theta(t)).$ (4.9)
Note that the $t\mapsto y(t)$ has been fixed until now. Thanks to the density result in Lemma A.5(), we know that there exists a sequence of definable selections $\Big(\underbrace{(x_{\ell},\lambda_{\ell},\mu_{\ell})}_{=:y_{\ell}}\Big)_{\ell\in\mathbb{N}}$ such that, for all $\ell\in\mathbb{N}$ and all $t$, $y_{\ell}(t)\in\mathrm{Sol_{pd}}(\theta(t))$ and such that, for all $t$,
$D_{f}(\theta(t))$ $=\{u\in\mathbb{R}^{q}:\exists\,y\in\mathrm{Sol_{pd}}(\theta(t)),(0,u)\in D_{\mathcal{L}}(y,\theta(t))\}$
$=\operatorname*{cl}\bigcup_{\ell\in\mathbb{N}}\{u\in\mathbb{R}^{q}:(0,u)\in D_{\mathcal{L}}(y_{\ell}(t),\theta(t))\}.$
For each $\ell\in\mathbb{N}$, $y_{\ell}$ satisfies (4.9) which applies to an arbitrary selection $y(t)=(x(t),\lambda(t),\mu(t))\in\mathrm{Sol_{pd}}(\theta(t))$. Using the same density argument, it follows, that, for almost every $t$,
$\frac{d}{dt}f(\theta(t))=\langle{\dot{\theta}(t),u}\rangle\quad\forall\,u\in D_{f}(\theta(t)).$
This concludes the proof. ∎
4.4 Failure of the method without the definability assumption
Let us illustrate the importance of the definability assumptions. For this we demonstrate that the local Lipschitz continuity, path differentiability and even differentiability of the objective function $F$, together with the continuous differentiability of the constraint functions $G$ and $H$, are not sufficient for $D_{f}$ in Theorem 4.6() to be a conservative field for $f$. The restriction to a subclass of functions with additional properties is required. As already stated, in this work, we elected for definable functions that arise in many relevant applications.
Before presenting the result, we describe a fractal-type set that will be central to the proof. Indeed, we define the set $C\subset\mathbb{R}^{2}$ as $C\coloneq\cap_{\ell\in\mathbb{N}}C_{\ell}$, where for each $\ell\in\mathbb{N}$, $C_{\ell}$ is the union of $4^{\ell}$ squares of side length $1/4^{\ell}$ each and $(C_{\ell})_{\ell\in\mathbb{N}}$ form a nested decreasing sequence for the inclusion. In addition, the set $C$ has the following properties:
  1. (i)
    $C$ is nonempty and closed with empty interior, and is included in $0,1]\times[0,1]$; hence it is compact;
  2. (ii)
    $C$ projected on the first axis gives $0,1]\times\{0\}$, and projected on the second axis it gives $\{0\}\times[0,1]$.
Such a set $C$ exists and can be constructed; see [34, Section 4.1] for all the details and all its properties. An illustration of the construction [34, Figure 1] is reproduced exactly in Figure 1().
Figure 1: The fractal construction in [34, Section 4.1]. They start with the closed unit square in black. It is divided into 16 squares, each with a side length equal to one fourth of the original square’s side length. Only four of them at specific positions are kept and the others dropped. This process is repeated recursively on each square ad infinitum. The additional red lines represent projection of these sets on rotated axes. Considering $C_{\ell}$, $\ell\in\mathbb{N}$, the set obtained after $\ell$ steps ($C_{0}$ is the original square). We have that $C_{\ell+1}\subset C_{\ell}$ for all $\ell$. Then, let $C=\cap_{\ell\in\mathbb{N}}C_{\ell}$, which is closed. The projection of $C$ on each axes are full segments. Furthermore, in the limit, both projections on rotated axes are Cantor sets of zero measure.
Claim 4.8.
There exists a locally Lipschitz objective function $F\colon\mathbb{R}^{2}\to\mathbb{R}$ that
  1. (i)
    is not definable in any o-minimal structure,
  2. (ii)
    is differentiable and path differentiable on $\mathbb{R}^{2}$, continuously differentiable on $\mathbb{R}^{2}\setminus C$, and
  3. (iii)
    is such that the set-valued map $D_{f}\colon\mathbb{R}\to\mathbb{R}$ given by
    $\theta$ $\mapsto\Big\{u\in\mathbb{R}:\exists\,(x,\lambda)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\partial^{c}F(x,\theta)+\left(\lambda\nabla_{\!x}G(x,\theta),0\right)\Big\}$
    is not a conservative field for the value function $f\colon\mathbb{R}\to\mathbb{R}$ defined by
    $\theta\mapsto\min_{x\in\mathbb{R}}\{F(x,\theta):G(x,\theta)\leq 0\},$
    where G is the continuously differentiable function $(x,\theta)\mapsto x^{2}-4$.
[…]
By Claim A.7(), there exists a differentiable, locally Lipschitz and path differentiable function $F\colon\mathbb{R}^{2}\to\mathbb{R}$ such that $F(x,\theta)=0$ for every $(x,\theta)\in C$, and $F(x,\theta)>0$ for every $(x,\theta)\in\mathbb{R}^{2}\setminus C$. In addition, $F$ is continuously differentiable everywhere except on $C$ with
$\mathcal{B}_{c}(0,1)\subset\partial^{c}F(x,\theta)\quad\forall\,(x,\theta)\in C.$ (4.10)
Given the fact that $\operatorname*{arg\,min}_{(x,\theta)\in\mathbb{R}^{2}}F(x,\theta)=C$ is not definable, $F$ is not definable either.
The set-valued map of feasible point $R\colon\mathbb{R}\to\mathbb{R}$, defined by
$\theta\mapsto\{x\in\mathbb{R}:G(x,\theta)\leq 0\},$
[…]
Since, $F$ is continuous and for every $\theta\in\mathbb{R}$, $R(\theta)=[-2,2]$, the set-valued map of primal solutions $\mathrm{Sol_{p}}\colon\mathbb{R}\to\mathbb{R}$, defined by
$\theta\mapsto\operatorname*{arg\,min}\{F(x,\theta):x\in R(\theta)\},$
[…]
Thanks to Definition $4\tagform@ii$(), for every $\theta\in[0,1]$, there exists $x\in[0,1]$ such that $(x,\theta)\in C$. It follows that the value function is zero on $[0,1]$, i.e,
$\forall\,\theta\in[0,1],\quad f(\theta)=0.$
Let $\theta\in[0,1]$ and let $\bar{x}\in\mathrm{Sol_{p}}(\theta)$. As $\bar{x}\in[0,1]$ and $G$ is therefore not active at $(\bar{x},\theta)$, we have $(\bar{x},0)\in\mathrm{Sol^{c}_{pd}}(\theta)$. It follows that
$D_{f}(\theta)$ $=\big\{u\in\mathbb{R}:\exists\,(x,\lambda)\in\mathrm{Sol^{c}_{pd}}(\theta),(0,u)\in\partial^{c}F(x,\theta)+\left(\lambda\nabla_{\!x}G(x,\theta),0\right)\big\}$
$\supset\big\{u\in\mathbb{R}:(0,u)\in\partial^{c}F(\bar{x},\theta)\big\}$
$\supset\big\{u\in\mathbb{R}:(0,u)\in\mathcal{B}_{c}(0,1)\big\}$
$=[-1,1],$
where we use (4.10) to get the second inclusion.
Consider the absolutely continuous curve $\gamma\colon[0,1]\to\mathbb{R}$ defined as $t\mapsto t$. Let’s now suppose, for the sake of contradiction, that $f$ has the chain rule property for $D_{f}$. Then, for almost all $t\in[0,1]$, and for any $u\in D_{f}(\gamma(t))$, in particular any $u\in[-1,1]$,
$\frac{d}{dt}f(\gamma(t))=\langle{u,\dot{\gamma}(t)}\rangle=\langle{u,1}\rangle=u.$
That’s impossible. So $f$ does not have the chain rule property for $D_{f}$, i.e, $D_{f}$ is not conservative for $f$. ∎
Appendix A Technical results
The following lemma is an immediate consequence of the stronger Aubin property result given in [32, Corollary 4.5]. We give its quick derivation in the proof just for completeness.
Lemma A.1.
The set-valued map of feasible points $R$ is inner semicontinuous, i.e., for all $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ and $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ such that $x^{*}\in R(\theta^{*})$ and $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$, there exists $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ such that $x_{k}\xrightarrow[k\to\infty]{}x^{*}$ and $x_{k}\in R(\theta_{k})$ for any $k\in\mathbb{N}$.
Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $x^{*}\in R(\theta^{*})$, and $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. Since $(x^{*},\theta^{*})$ is MFCQ-qualified, [32] proved that $R$ has the Aubin property, or equivalently, is pseudo-Lipschitzian, around $(x^{*},\theta^{*})$. That means there exist a neighborhood $\mathcal{N}$ of $x^{*}$, a neighborhood $\mathcal{U}$ of $\theta^{*}$ and $M\geq 0$ such that
$R(\theta_{1})\cap\mathcal{N}\subset R(\theta_{2})+M\lVert{\theta_{1}-\theta_{2}}\rVert\mathcal{B}_{c}(0,1)\qquad\forall\,\theta_{1},\theta_{2}\in\mathcal{U}.$
In particular, for any $k\geq K$, with $K$ large enough,
$\{x^{*}\}\subset R(\theta^{*})\cap\mathcal{N}\subset R(\theta_{k})+M\lVert{\theta^{*}-\theta_{k}}\rVert\mathcal{B}_{c}(0,1).$
Therefore, for any $k\geq K$ , there exists $x_{k}\in R(\theta_{k})$ and $z_{k}\in\mathcal{B}_{c}(0,1)$ such that
$x^{*}=x_{k}+M\lVert{\theta^{*}-\theta_{k}}\rVert z_{k}.$
It follows that
$\lVert{x^{*}-x_{k}}\rVert\leq M\lVert{\theta^{*}-\theta_{k}}\rVert\text{ and }\lim_{k\to\infty}\lVert{x^{*}-x_{k}}\rVert=0.$
So $x_{k}\xrightarrow[k\to\infty]{}x^{*}$, which proves that $R$ is inner semicontinuous. ∎
Lemma A.2.
The set-valued map of primal solutions $\mathrm{Sol_{p}}$ has a closed graph.
Let $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$, $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ and $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})$ for all $k\in\mathbb{N}$, $x_{k}\xrightarrow[k\to\infty]{}x^{*}$ and $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. It’s immediate that $x^{*}\in R(\theta^{*})$.
Take $y\in R(\theta^{*})$. Since $R$ is inner semicontinuous thanks to Lemma A.1(), there exists $(y_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ such that $y_{k}\xrightarrow[k\to\infty]{}y$ and $y_{k}\in R(\theta_{k})$ for any $k\in\mathbb{N}$. It follows that, for every $k\in\mathbb{N}$,
$F(x_{k},\theta_{k})\leq F(y_{k},\theta_{k}),$
which implies that
$\lim_{k\to\infty}F(x_{k},\theta_{k})\leq\lim_{k\to\infty}F(y_{k},\theta_{k}),$
and, by continuity of $F$,
$F(x^{*},\theta^{*})\leq F(y,\theta^{*}).$
We can conclude that $x^{*}\in\mathrm{Sol_{p}}(\theta^{*})$ and that $\mathrm{Sol_{p}}$ is graph closed. ∎ The following lemma can also be obtained through the Berge’s maximum theorem.
Lemma A.3.
The value function $f$, defined in (1.1), is continuous on $\mathbb{R}^{q}$.
Let $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ and $\theta^{*}\in\mathbb{R}^{q}$ such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. Take $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})$ for each $k\in\mathbb{N}$. Since $\mathrm{Sol_{p}}$ is locally bounded, there exists a neighborhood $\mathcal{N}$ of $\theta^{*}$ and $r>0$ such that
$\bigcup_{\theta\in\mathcal{N}}\mathrm{Sol_{p}}(\theta)\subset\mathcal{B}_{c}(0,r).$
For $K$ large enough, we have $\theta_{k}\in\mathcal{N}$ and $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})\subset\mathcal{B}_{c}(0,r)$ whenever $k\geq K$. We then get that $(x_{k})_{k\in\mathbb{N}}$ is bounded. Therefore $(f(\theta_{k}))_{k\in\mathbb{N}}=(F(x_{k},\theta_{k}))_{k\in\mathbb{N}}$ is bounded by continuity of $F$.
Let $f^{*}\in\mathbb{R}$ be a limit point of $(f(\theta_{k}))_{k\in\mathbb{N}}$. Thus there exists $(f(\theta_{k_{s}}))_{s\in\mathbb{N}}$ such that $f(\theta_{k_{s}})\xrightarrow[s\to\infty]{}f^{*}$. The sequence $(\theta_{k_{s}})_{s\in\mathbb{N}}$, being a subsequence of $(\theta_{k})_{k\in\mathbb{N}}$, converges to $\theta^{*}$. The corresponding sequence $(x_{k_{s}})_{s\in\mathbb{N}}$ is bounded by the previous paragraph. So, we can extract a convergent subsequence $(x_{k_{s_{\ell}}})_{\ell\in\mathbb{N}}$ and let $x^{*}$ be its limit. Since $\mathrm{Sol_{p}}$ has a closed graph by Lemma A.2(), $x^{*}\in\mathrm{Sol_{p}}(\theta^{*})$. It follows that
$f^{*}=\lim_{s\to\infty}f(\theta_{k_{s}})=\lim_{\ell\to\infty}f(\theta_{k_{s_{\ell}}})=\lim_{\ell\to\infty}F(x_{k_{s_{\ell}}},\theta_{k_{s_{\ell}}})=F(x^{*},\theta^{*})=f(\theta^{*}).$
As $f^{*}$ was arbitrary, we get that $f(\theta^{*})$ is the only accumulation point of the bounded sequence $(f(\theta_{k}))_{k\in\mathbb{N}}$. Consequently, $f(\theta_{k})\xrightarrow[k\to\infty]{}f(\theta^{*})$ and $f$ is continuous. ∎
The following two lemmas are definable versions of Castaing representation theorem.
Lemma A.4 ([34, Lemma 8]).
Let $J\colon\mathbb{R}^{s}\rightrightarrows\mathbb{R}^{r}$ be a compact-valued definable set-valued map with nonempty values. Then there exists a sequence of definable selections $(V_{\ell})_{{\ell}\in\mathbb{N}}$ for $J$ such that, for any $x\in\mathbb{R}^{s}$,
$J(x)=\operatorname*{cl}\{V_{\ell}(x):\ell\in\mathbb{N}\}.$
Lemma A.5 ([34, Lemma 9]).
Let Assumptions 4.2() hold. Let $D_{F}$ a be a definable conservative field for $F$. Then there exists a sequence of definable selections $\left\{(x_{\ell},\lambda_{\ell},\mu_{\ell})\right\}_{{\ell}\in\mathbb{N}}$ such that, for any $\ell$$\in\mathbb{N}$ and $\theta\in\mathbb{R}^{q}$, $(x_{\ell}(\theta),\lambda_{\ell}(\theta),\mu_{\ell}(\theta))\in\mathrm{Sol_{pd}}(\theta)$ and such that, for all $\theta\in\mathbb{R}^{q}$,
$\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\operatorname*{conv\!}D_{F}(x,\theta)$
$\qquad\quad+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right)\Big\}$
$=\operatorname*{cl}\big\{u\in\mathbb{R}^{q}:\exists\,{\ell}\in\mathbb{N},(0,u)\in\operatorname*{conv\!}D_{F}(x_{\ell}(\theta),\theta)$
$\qquad\quad+\big(\operatorname{Jac}_{x}\!G(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\lambda_{\ell}(\theta)+\operatorname{Jac}_{x}\!H(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\mu_{\ell}(\theta),\operatorname{Jac}_{\theta}\!G(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\lambda_{\ell}(\theta)$
$\qquad\qquad\qquad+\operatorname{Jac}_{\theta}\!H(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\mu_{\ell}(\theta)\big)\big\}.$
Lemma A.6.
Let $A\subset\mathbb{R}^{M}$ be closed with empty interior. Then $A$ is the set of accumulation points of a sequence $(z_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{M}\setminus A$ with $z_{s}\neq z_{r}$ for all $s\neq r$.
Since every subset of $\mathbb{R}^{M}$ is itself separable, there exists a sequence $(c_{r})_{r\in\mathbb{N}}\subset A$ which is dense in $A$. We know that $\mathbb{R}^{M}\setminus A$ is dense in $\mathbb{R}^{M}$. Therefore, for every $\ell\in\mathbb{N}$ and $s\in\{1,\ldots,\ell\}$, we can take
$w_{\ell,s}\in(\mathbb{R}^{M}\setminus A)\cap\mathcal{B}(c_{s},1/\ell)$
such that
$w_{\ell,s}\notin\{w_{r,t}:r,t\in\mathbb{N},r<\ell,\text{ and }1\leq t\leq r\}\cup\{w_{r,t}:r,t\in\mathbb{N},r=\ell,\text{ and }1\leq t<s\}.$
We set $(z_{k})_{k\in\mathbb{N}}$ as an ordered enumeration of $\{w_{1,1},w_{2,1},w_{2,2},w_{3,1},w_{3,2},w_{3,3},w_{4,1},\ldots\}.$
Let $\bar{z}\in A$. Then there exists $(c_{r_{s}})_{s\in\mathbb{N}}\subset A$ such $c_{r_{s}}\xrightarrow[s\to\infty]{}\bar{z}$. Let $\varepsilon>0$ and $N\in\mathbb{N}$. There exist $s\in\mathbb{N}$ such that $\lVert{{c_{r_{s}}}-{\bar{z}}}\rVert<\varepsilon/2$. Let $L\in\mathbb{N}$ such that $1/L<\varepsilon/2$ and $L>r_{s}$. We have $w_{\ell,r_{s}}\in\mathcal{B}(c_{r_{s}},1/\ell)$ for all $\ell\geq L$. It follows that $\lVert{{w_{\ell,r_{s}}}-{\bar{z}}}\rVert<\varepsilon$ for all $\ell\geq L$. Let $\ell\geq\max\{L,N\}$. By definition, there exists $k\geq N$ such that $z_{k}=w_{\ell,r_{s}}$. Then $\lVert{{z_{k}}-{\bar{z}}}\rVert<\varepsilon$ and $\bar{z}$ is an accumulation point of $(z_{k})_{k\in\mathbb{N}}$.
Let $\bar{z}\not\in A$. Since $A$ is closed, we have that $\operatorname*{dist}(\bar{z},A)>0$, with
$\operatorname*{dist}(\bar{z},A)=\inf\{\lVert{{\bar{z}}-{a}}\rVert:a\in A\}.$
Let $L\in\mathbb{N}$ such that $1/L<\operatorname*{dist}(\bar{z},A)/2$. So for every $\ell\geq L$, for all $s\in\{1,\ldots,\ell\}$, $\lVert{{c_{s}}-{w_{\ell,s}}}\rVert<1/\ell\leq 1/L$ . Suppose that there exist $\ell\geq L$ and $s\in\{1,\ldots,\ell\}$ such that $\lVert{{w_{\ell,s}}-{\bar{z}}}\rVert<1/L$. It follows that
$2/L<\operatorname*{dist}(\bar{z},A)\leq\lVert{{\bar{z}}-{c_{s}}}\rVert\leq\lVert{{\bar{z}}-{w_{\ell,s}}}\rVert+\lVert{{w_{\ell,s}}-{c_{s}}}\rVert<1/L+1/L=2/L.$
This is a contradiction. So for every $\ell\geq L$, $\lVert{{\bar{z}}-{w_{\ell,s}}}\rVert\geq 1/L$ for all $s\in\{1,\ldots,\ell\}$. That means that there exists $K\in\mathbb{N}$ such that for any $k\geq K$, $\lVert{{z_{k}}-{\bar{z}}}\rVert\geq 1/L$. Thus $\bar{z}$ is not an accumulation point of $(z_{k})_{k\in\mathbb{N}}$.
We can conclude that the accumulation points of $(z_{k})_{k\in\mathbb{N}}$ are exactly $A$. ∎
Claim A.7.
Let $C\subset\mathbb{R}^{\ell}$ be closed with empty interior. Then, there exists a differentiable and locally Lipschitz function $\bar{g}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ such that $\bar{g}(z)=0$ for every $z\in C$, and $\bar{g}(z)>0$ for every $z\in\mathbb{R}^{\ell}\setminus C$. Moreover $\bar{g}$ is continuously differentiable everywhere except on $C$ with $\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z)$ for any $z\in C$. In particular, if $C$ is defined as in Section 4.4(), $\bar{g}$ is also path differentiable.
Consider the bump function $\bar{b}\colon\mathbb{R}\to\mathbb{R}$ defined by
$\bar{b}(x)\coloneq\begin{cases}e^{-1/(1-x^{2})}&\text{if }|x|<1,\\ 0&\text{if }|x|\geq 1.\end{cases}$
Let $b(x)=\left(\bar{b}(x)\right)^{2}$. The derivative $b^{{}^{\prime}}$ is odd and $b^{{}^{\prime}}(1/2)<0$. Define $\varphi\colon\mathbb{R}\to\mathbb{R}$ as $\varphi(x)=\frac{1}{-b^{{}^{\prime}}(1/2)}b(x)$. The function $\varphi$ is infinitely differentiable and is support is [1,1][-1,1]. In addition, $\varphi^{\prime}(1/2)=-1$ and $\varphi^{\prime}(-1/2)=-\varphi^{\prime}(1/2)=1$.
By Whitney theorem [30, Theorem 2.29], $C$ is the zero of an infinitely differentiable and non-negative function $\bar{f}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ and, for any $z\in C$, $\nabla\bar{f}(z)=0$. Thanks to Lemma A.6(), there exists a sequence $(z_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{2}\setminus C$ such that its set of accumulation points is $C$. For any $k\in\mathbb{N}$, consider $\mathcal{B}(z_{k},r_{k})$ such that $r_{k}>0$, $r_{k}\xrightarrow[k\to\infty]{}0$, and $\mathcal{B}(z_{k},r_{k})\cap\mathcal{B}(z_{s},r_{s})=\varnothing$ for $s\neq k$. Such a choice is possible, because for each $k$, $z_{k}$ is at positive distance of $z_{i}$ for each $i\neq k$ and at positive distance of $C$ the set of accumulation points. Define $\bar{g}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ as
$z\mapsto\bar{f}(z)+\sum_{k\in\mathbb{N}}\varphi\left(\frac{\lVert{z-z_{k}}\rVert}{r_{k}}\right)r_{k}.$
Then, $\bar{g}(z)=0$ for every $z\in C$, and $\bar{g}(z)>0$ for every $z\in\mathbb{R}^{\ell}\setminus C$. Moreover $\bar{g}$ is differentiable and locally Lipschitz on $\mathbb{R}^{\ell}$. It is continuously differentiable everywhere except on $C$ because $\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z)$ for any $z\in C$. Indeed, let $\bar{z}\in C$ and, without loss of generality, we assume that $z_{k}\xrightarrow[k\to\infty]{}\bar{z}$. Let $v\in\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}$. Consider $(\bar{z}_{k})_{k\in\mathbb{N}}$ such that $\bar{z}_{k}=z_{k}-\frac{r_{k}}{2}v$ for all $k\in\mathbb{N}$. Then we have, for any $k\in\mathbb{N}$, $\nabla\bar{g}(\bar{z}_{k})=\nabla\bar{f}(\bar{z_{k}})+v$. So, $\bar{z}_{k}\xrightarrow[k\to\infty]{}\bar{z}$ and $\nabla\bar{g}(\bar{z}_{k})\xrightarrow[k\to\infty]{}v$. It follows that $v\in\partial^{c}\bar{g}(\bar{z})$. Since $v$ is arbitrary in $\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}$, and by Definition 3.1(), we have
$\operatorname*{conv\!}\,\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}=\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z).$
Now let us assume that $C$ is defined as in Section 4.4(). Let $\gamma\colon[0,1]\to\mathbb{R}^{\ell}$ be an absolutely continuous curve. Given that $\bar{g}$ is locally Lipschitz, $F\circ\gamma$ is absolutely continuous. Let $\Omega\subset[0,1]$ be the full measure set where $F\circ\gamma$ and $\gamma$ are differentiable. Consider the following three sets:
$S=\{t\in\Omega:\gamma(t)\in\mathbb{R}^{\ell}\setminus C\},$
$S_{1}=\{t\in\Omega:\gamma(t)\in C\text{ and }\dot{\gamma}(t)=0\},$
$S_{2}=\{t\in\Omega:\gamma(t)\in C\text{ and }\dot{\gamma}(t)\neq 0\}.$
We have $\Omega=S\cup S_{1}\cup S_{2}$. Since $\bar{g}$ is continuously differentiable outside $C$, the chain rule holds on $S$, i.e., for any $t\in S$,
$\frac{d}{dt}\bar{g}(\gamma(t))=\langle{u,\dot{\gamma}(t)}\rangle\quad\forall\,u\in\partial^{c}\bar{g}(\gamma(t))=\{\nabla\bar{g}(\gamma(t))\}.$
Thanks to [34, Lemma 6], we know that $S_{2}$ has zero measure. As $\bar{g}$ is locally Lipschitz and $\gamma(t+h)=\gamma(t)+o(h)$ for all $(t+h,t)\in\Omega\times S_{1}$ with $|h|$ small enough, the chain rule holds in $S_{1}$. In conclusion, the chain rule holds for almost all $t\in[0,1]$ and $\bar{g}$ is path differentiable. ∎
Acknowledgements.
The authors warmly thank Bruno Després and François Pacaud for relevant discussions and suggestions.
This work has been supported by the Occitanie region, the European Regional Development Fund (ERDF), and the French government, through the France 2030 project managed by the National Research Agency (ANR) with the reference number “ANR-22-EXES-0015.
JB and EP thank AI Interdisciplinary Institute ANITI funding, through the ANR under the France 2030 program (grant ANR-23-IACL-0002), Chair TRIAL, Air Force Office of Scientific Research, Air Force Material Command, USAF, under grant numbers FA8655-22-1-7012. JB and EP acknowledge support from ANR MAD. JB and EP thank TSE-P and acknowledge support from ANR Chess, grant ANR-17-EURE-0010, ANR Regulia. EP acknowledges support from IUF. As part of the Madlearning project, this work received funding from the French government, managed by the National Research Agency (ANR) under the France 2030 program, with reference number ’ANR-25-PEIA-0002’.